OpenStudy (anonymous):

solve by substitution -x^2 - x + 19 = y x=y+80

OpenStudy (anonymous):

@jim_thompson5910 could you help

OpenStudy (anonymous):

@amistre64

jimthompson5910 (jim_thompson5910):

first solve x=y+80 for y and tell me what you get

OpenStudy (anonymous):

@jim_thompson5910 you can't because both variables are unknown, right?

jimthompson5910 (jim_thompson5910):

you can still isolate y though

jimthompson5910 (jim_thompson5910):

get y all by itself on one side

OpenStudy (anonymous):

so it would be x-80=y ??

jimthompson5910 (jim_thompson5910):

good, y = x-80

jimthompson5910 (jim_thompson5910):

now we turn to -x^2 - x + 19 = y

jimthompson5910 (jim_thompson5910):

and replace very copy of y in -x^2 - x + 19 = y with x-80 (since we found that y = x - 80)

jimthompson5910 (jim_thompson5910):

that gives us -x^2 - x + 19 = y -x^2 - x + 19 = x-80

jimthompson5910 (jim_thompson5910):

do you know how to take it from here?

OpenStudy (anonymous):

@jim_thompson5910 no i'm confused :(

jimthompson5910 (jim_thompson5910):

but do you see how I got -x^2 - x + 19 = x-80 ?

OpenStudy (anonymous):

@jim_thompson5910 yes, you substituted y with x-80 because y= x-80

jimthompson5910 (jim_thompson5910):

now we need to get everything to one side

jimthompson5910 (jim_thompson5910):

-x^2 - x + 19 = x-80 -x^2 - x + 19 - x = -80 -x^2 - x + 19 - x + 80 = 0 -x^2 - 2x + 99 = 0 what's the next step?

OpenStudy (anonymous):

@jim_thompson5910 isolate x ?? so add 99 to both sides?

jimthompson5910 (jim_thompson5910):

not quite, that would work if we didn't have that x^2 term in there

jimthompson5910 (jim_thompson5910):

but we can do this multiply everything by -1 to make the x^2 term positive -x^2 - 2x + 99 = 0 x^2 + 2x - 99 = 0 basically the signs of each term change

jimthompson5910 (jim_thompson5910):

Now factor x^2 + 2x - 99 you need to find two numbers that multiply to -99 and add to 2

OpenStudy (anonymous):

@jim_thompson5910 -99 isn't divisible by anything though?

jimthompson5910 (jim_thompson5910):

99 = 11*9 so -99 = 11*(-9) or -99 = -11*9

OpenStudy (anonymous):

@jim_thompson5910 oh okay. so 11 and -9 ?

jimthompson5910 (jim_thompson5910):

good, so that means x^2 + 2x - 99 = 0 turns into (x + 11)(x - 9) = 0

jimthompson5910 (jim_thompson5910):

if (x + 11)(x - 9) = 0 then x = ??? or x = ???

OpenStudy (anonymous):

@jim_thompson5910 okay.. so then it would be x= 9 and x= -11?

jimthompson5910 (jim_thompson5910):

good

jimthompson5910 (jim_thompson5910):

once you know that, you can find y

jimthompson5910 (jim_thompson5910):

if x = 9, then y is... y = x - 80 y = 9 - 80 ... replace x with 9 since x = 9 y = -71 one ordered pair is (9, -71)

jimthompson5910 (jim_thompson5910):

what's the other ordered pair?

OpenStudy (anonymous):

@jim_thompson5910 okay one sec

OpenStudy (anonymous):

@jim_thompson5910 -91??

OpenStudy (anonymous):

so -11, -91?

jimthompson5910 (jim_thompson5910):

(-11, -91) don't forget the parenthesis

jimthompson5910 (jim_thompson5910):

you are correct though

jimthompson5910 (jim_thompson5910):

these two ordered pairs represent the points of intersection between the two graphs

OpenStudy (anonymous):

@jim_thompson5910 YAY THANK YOU SO MUCH <3

jimthompson5910 (jim_thompson5910):

you're welcome